OpenStudy (babynini):

Solve each equation in the interval [0,2pi) rounded to two decimal points a) (2cos(theta)-1)(sin(theta)-1)=0

OpenStudy (babynini):

em I got 2cos(theta)sin(theta)-2cos(theta)-sin(theta)+1=0 yeah?

OpenStudy (freckles):

If a*b=0, then either a=0 or b=0 or both=0. This means you have: \[2\cos(\theta)-1=0 \text{ or } \sin(\theta)-1=0\]

OpenStudy (freckles):

isolate the trig function

OpenStudy (freckles):

\[\cos(\theta)=\frac{1}{2} \text{ or } \sin(\theta)=1 \]

OpenStudy (freckles):

try to solve both of those equations

OpenStudy (babynini):

cos(theta) = 1.047 sin(theta) = 1.57

OpenStudy (freckles):

oh you mean: \[\theta \approx 1.05 \text{ you should also be able to obtain } \theta \approx 5.24 \text{ from } \cos(\theta)=\frac{1}{2}\] and \[\theta \approx 1.57 \text{ from the other equation } \sin(\theta)=1 \]

OpenStudy (babynini):

Yes, that is what I meant :) how did you get 5.24?

OpenStudy (freckles):

\[\cos(\theta)=\cos(-\theta)=\cos(-\theta+2 \pi ) \\ \cos(1.05) \approx \cos(-1.05+2\pi) \approx \cos(-1.05+2(3.14)) \\ \cos(1.05) \approx \cos(-1.05+6.28)=\cos(5.23) \\ \text{ but honestly I used the unit circle }\]

OpenStudy (freckles):

You see the points on the circle. The first number of every pair is the output for the cos(the angle there). The second number of every pair is the output for the sin(the angle there). I was looking for when the x-coordinate (the first number aka the cos number) was 1/2. I see this was happening at the pi/3 mark and the 5pi/3 mark.

OpenStudy (freckles):

And your question wanted these values rounded to the nearest hundredth.

OpenStudy (freckles):

\[\frac{\pi}{3} \approx 1.05 \\ \frac{5 \pi}{3} \approx 5.24\]

OpenStudy (babynini):

ooo I see! that makes sense! haha why do we not want to do the othe pi/3 s?

OpenStudy (freckles):

at 2pi/3 and 4pi/3 you should see the cos number (the x-coordinate) is negative

OpenStudy (freckles):

1/2 is not a negative number

OpenStudy (babynini):

at pi/3 it is not negative either. That was another question I have why do we make it -1.05 instead of adding 2pi to 1.05

OpenStudy (freckles):

cos is an even function \[\cos(\theta)=\cos(-\theta)\]

OpenStudy (freckles):

sin is an odd function \[\sin(\theta)=-\sin(-\theta)\]

OpenStudy (babynini):

ooh yes.

OpenStudy (freckles):

and then inside you can add as many 2pi as you like because we just wind up at the same place

OpenStudy (freckles):

but I add one 2pi since that would give me another solution in [0,2pi)

OpenStudy (freckles):

Pretend we want to solve this: \[\cos(\theta)=.23 , \theta \in [0,2\pi)\] This .23 is not on the unit circle. But using the unit circle and other identities helps me to solve this equation. First I can do arccos( ) on both sides to find one solution. \[\theta=\arccos(.23) \] another solution would be given by: \[-\theta+2\pi=\arccos(.23) \\ \theta=-\arccos(.23)+2\pi\] and if we wanted to approximate these solutions by rounding to the nearest hundredth we would have: \[\theta \approx 4.94 , 1.34 \]

OpenStudy (freckles):

The second solution I used the fact that cos was even and had a period of 2pi.

OpenStudy (babynini):

Hmm ok.

OpenStudy (babynini):

http://openstudy.com/users/babynini#/updates/55823438e4b07028ea611c18 so for this one, how do I do that?

OpenStudy (babynini):

Thank you for the help on this one, btw. :)

OpenStudy (freckles):

np